OpenStudy (anonymous):

(I have to pass) Determine which ordered pairs, if any, are solutions to the system: y=x^2-6x+5 y=2x-7 a.(2,-3) b.(1,-5) c.(6,5)

OpenStudy (anonymous):

ANSWERS ARE! Choices a and c both are solutions to the system Choice a is the only ordered pair given that is a solution to the system None of these ordered pairs are solutions Choices a and b both are solutions to the system

OpenStudy (anonymous):

lol i have to pass! you need the answers?

OpenStudy (anonymous):

I got B.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3Dx^2-6x%2B5%2Cy%3D2x-7 go with Choices a and c both are solutions to the system

OpenStudy (anonymous):

It's A??

OpenStudy (solomonzelman):

y=x^2-6x+5 y=x^2-6x+9-9+5 y=(x-3)^2²-4 compare to y=a(x-h)²+k then go from there.

OpenStudy (anonymous):

I have to be sure the answer is certain..